OpenStudy (cutiecomittee123):

HELP TRIG QUESTIONS a. One solution to sintheta=sqrt3/2 is theta= b. One angle that is coterminal to the angle you found in part a is theta= c. Another solution to the equation sintheta=sqrt3/2 is theta= d. One angle that is coterminal to the angle that you found in part c is theta=

OpenStudy (anonymous):

sorry i aint there yet

OpenStudy (cutiecomittee123):

darn

jimthompson5910 (jim_thompson5910):

use the unit circle http://etc.usf.edu/clipart/43200/43215/unit-circle7_43215_lg.gif find all of the points that have a y coordinate of sqrt(3)/2, then find the corresponding value of theta

OpenStudy (cutiecomittee123):

there are only two points that have the y coordinate sqrt(3)/2 which are 2pi/3 and pi/3

OpenStudy (cutiecomittee123):

and those two are coterminal to eachother

OpenStudy (cutiecomittee123):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

2pi/3 and pi/3 aren't coterminal

jimthompson5910 (jim_thompson5910):

but you have the right values of theta

OpenStudy (cutiecomittee123):

so confused then

OpenStudy (anonymous):

which one you working on?

OpenStudy (anonymous):

\[\sin(\theta)=\frac{\sqrt3}{2}\]?

OpenStudy (cutiecomittee123):

yes

OpenStudy (anonymous):

find a place on the unit circle where the second coordinate is \(\frac{\sqrt3}{2}\) there are a couple of them the angle is your answer

jimthompson5910 (jim_thompson5910):

Step 1) find all points that have a y coordinate of sqrt(3)/2 step 2) Find the corresponding angles to those points. Those angles are pi/3 and 2pi/3 So that's why \(\Large \theta = \frac{\pi}{3}\) is a solution to \(\Large \sin(\theta)=\frac{\sqrt3}{2}\) Also, \(\Large \theta = \frac{2\pi}{3}\) is another solution to \(\Large \sin(\theta)=\frac{\sqrt3}{2}\)

jimthompson5910 (jim_thompson5910):

Stated another way: \(\Large \sin\left(\frac{\pi}{3}\right)=\frac{\sqrt3}{2}\) \(\Large \sin\left(\frac{2\pi}{3}\right)=\frac{\sqrt3}{2}\)

OpenStudy (cutiecomittee123):

so then i was right.

OpenStudy (cutiecomittee123):

but if there are only two and c. and d. are asking for the same thing then huh? how can there be four of them

jimthompson5910 (jim_thompson5910):

`b. One angle that is coterminal to the angle you found in part a is theta=` simply add 2pi to the answer you stated in part a)

OpenStudy (cutiecomittee123):

4pi/3 and 2pi/3 again

jimthompson5910 (jim_thompson5910):

so if you have pi/3 for part a) then pi/3 + 2pi = pi/3 + 2pi*(3/3) pi/3 + 2pi = pi/3 + 6pi/3 pi/3 + 2pi = 7pi/3 the angles pi/3 and 7pi/3 are coterminal angles

OpenStudy (cutiecomittee123):

Oh i see

jimthompson5910 (jim_thompson5910):

if you have 2pi/3 for part a) then 2pi/3 + 2pi = 2pi/3 + 2pi*(3/3) 2pi/3 + 2pi = 2pi/3 + 6pi/3 2pi/3 + 2pi = 8pi/3 the angles 2pi/3 and 8pi/3 are coterminal angles

jimthompson5910 (jim_thompson5910):

`c. Another solution to the equation sintheta=sqrt3/2 is theta= ` simply state the other solution that wasn't mentioned in part a)

jimthompson5910 (jim_thompson5910):

`d. One angle that is coterminal to the angle that you found in part c is theta=` add 2pi to whatever your answer for part c) is

OpenStudy (cutiecomittee123):

okay i think i get that